OpenStudy (anonymous):

\[f(x) =\begin{cases}\frac{xy}{\sqrt{x^2+y^2}} & (x,y)\ne 0 \\0 & (x,y) = 0\end{cases}\] a) Determine if f is differentiable at (0,0) b) determine for which unit vectors v = vi + wj the directional derivative \(D_vf(0, 0)\) exists.

OpenStudy (anonymous):

\[h(0,0) = f(0,0) + f_x(0,0)(x-0)+f_y(0,0)(y-0) = 0\] \[\lim_{(x,y) \rightarrow(0,0)}\frac{f(x,y)-h(x,y)}{||(x,y)-(0,0)||} \] What I've got is limit does not exist, so it is not differentiable. Is it correct?

OpenStudy (anonymous):

Differentiable with respect to all directions?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

I'm not sure what definition you're using for directional derivative...

OpenStudy (anonymous):

\[D_vf(a) = \lim_{h\rightarrow 0}\frac{f(a+hv) - f(a)}{h}\]

OpenStudy (anonymous):

\[D_{\langle a,b\rangle }f(x, y) = \lim_{h\rightarrow 0}\frac{f(x+ah, y+bh) - f(x,y)}{h}\]

OpenStudy (anonymous):

Oh wait, so it is a single variable function then?

OpenStudy (anonymous):

a and v are vectors.

OpenStudy (anonymous):

\[ \lim_{h\to 0}\frac{f(ah,bh)-f(0,0)}{h} \]

OpenStudy (anonymous):

Are you talking about (a) or (b)?

OpenStudy (anonymous):

I'm saying \(v=\langle a,b \rangle\)

OpenStudy (anonymous):

I'm starting with a)

OpenStudy (anonymous):

\[\Huge \lim_{h\to 0} \frac{\frac{abh^2}{\sqrt{a^2h^2+b^2h^2}}}h \]

OpenStudy (anonymous):

That gives\[\frac{ab}{\sqrt{a^2+b^2}}\]

OpenStudy (anonymous):

\[ a^2+b^2>0 \]

OpenStudy (anonymous):

The \(\mathbf 0\) vector is only direction that it wouldn't be defined I guess? But technically that sort of vector isn't to be used for a directional derivative.

OpenStudy (anonymous):

I think so, or else, we'll have a new function for this question. But isn't the definition of differentiability of f at (a,b) like this: \[h(x,y) = f(a,b) + f_x(a,b)(x-a)+f_y(a,b)(y-b)\] \[\lim_{(x,y)\rightarrow(a,b)} \frac{f(x,y)-h(x,y)}{||(x,y)-(a,b)||}=0\] if f is differentiable at (a,b)?

OpenStudy (anonymous):

I got mine from http://tutorial.math.lamar.edu/Classes/CalcIII/DirectionalDeriv.aspx

OpenStudy (anonymous):

I've always learned that differentiablility just means that the limit of the difference quotient exists at the point in question.

OpenStudy (anonymous):

The first one isn't asking for directional derivative?

OpenStudy (anonymous):

Hmm, I see, I'm not sure how differentiability would be defined here.

OpenStudy (anonymous):

So it would be that there is a unique tangent plane then..?

OpenStudy (anonymous):

I think I understand your definition now. So what is the problem?

OpenStudy (anonymous):

Just to confirm if the limit doesn't limit.

OpenStudy (anonymous):

for part a

OpenStudy (anonymous):

I'm not so sure. You're getting a \(0/0\) which is an indeterminant form?

OpenStudy (anonymous):

\[\huge\lim_{(x,y) \rightarrow(0,0)}\frac{\frac{xy}{\sqrt{x^2+y^2}}}{\sqrt{(x-0)^2+(y-0)^2}} \]

OpenStudy (anonymous):

Is this what you went through?

OpenStudy (anonymous):

\[ \lim_{(x,y)\to (0,0)}\frac{xy}{x^2+y^2} \]

OpenStudy (anonymous):

Then you realized this limit doesn't exist?

OpenStudy (anonymous):

Maybe I'm retracing your steps here. Since the problem isn't obvious I have to work it out as well.

OpenStudy (anonymous):

From that, change it to \[\lim_{r \rightarrow 0} sin\theta cos\theta \ne 0\]So, it is not differentiable at (0,0)

OpenStudy (anonymous):

Yes, that looks right.

OpenStudy (anonymous):

Wow! *relieved*

OpenStudy (anonymous):

Different thetas (directions paths of the limit) are giving different values. Not a good sign for a limit.

OpenStudy (anonymous):

They don't necessarily have to be 0, they have to be constant.

OpenStudy (anonymous):

No, it gives different values, then limit doesn't exist. When limit exists, it should approach to a certain value...

OpenStudy (anonymous):

And from my textbook, it should be 0... o_o

OpenStudy (anonymous):

I mean it should be 0 for it to be differentiable.

OpenStudy (anonymous):

Nope. Unless they are using some other premise to suggest it couldn't be constant as some non 0 value and then asserting it'd have to be 0

OpenStudy (anonymous):

Here's the definition from the book

OpenStudy (anonymous):

Oh, the book said it had to be \(=0\). I didn't know that yet.

OpenStudy (anonymous):

I was thinking the limit just had to exist

OpenStudy (blockcolder):

Has it been shown that \[D_{\mathbb{u}} f(x,y)=\nabla{f}\cdot \mathbb{u}?\]

OpenStudy (anonymous):

It has been mentioned in the book that \[D_{v}f(a) = \nabla f(a)\cdot v\]

OpenStudy (anonymous):

That is useful when assuming a function is differentiable.

OpenStudy (anonymous):

Hmm.. But it is not differentiable in this case?

OpenStudy (anonymous):

I mean differentiable in terms of the directional derivative, not in terms of that weird definition we used in a)

OpenStudy (anonymous):

Honestly I'm not completely sure what the use for that definition is exactly.

OpenStudy (anonymous):

We should use this

OpenStudy (anonymous):

I already did it.

OpenStudy (anonymous):

Hmm, I guess we are done with this question! Thanks a lot!!

OpenStudy (anonymous):

Yeah.